sam (.sam.):

See if you guys can integrate this :) ∫3/(x^2+2) dx

sam (.sam.):

\[\int\limits_{}^{} \frac{3}{x ^{2}+2} dx\]

OpenStudy (ash2326):

\[\int \frac{3}{x^2+2} dx\] \[x= \sqrt {2} \tan \theta\] \[dx=\sqrt{2} \sec^2 \theta\] \[\int3 \sqrt{2}\sec^2 \theta \frac{d\theta }{2\tan^2\theta+2}\] \[\int 3\sqrt{2} \sec^2 \theta \frac{d\theta}{2 \sec^2\theta}\] we get \[\int 3 \sqrt{2} \frac{d\theta}{2}\] we get on integration \[ \frac{3}{\sqrt{2}} \theta+C\] we have \[ \tan \theta= \sqrt{2} x\] \[\theta = \tan ^{-1} (\sqrt {2} x)\] so we get \[ \frac{3}{\sqrt{2}}\tan ^{-1} (\sqrt {2} x)+ C \]

OpenStudy (ash2326):

we have x^2+2 in the denominator if \[x= \sqrt {2} x\] then \[ x^2+2= 2 \tan^2 \theta +2= 2(\tan^2 \theta +1)= 2 \sec^2 \theta\] also \[dx =\sqrt {2} \sec^2 \theta\] so the \(\sec^2 \theta \) gets cancelled from numerator and denominator, we 're left with an easy integration.

OpenStudy (anonymous):

hey sam can you help? x^2+18x+4=0

OpenStudy (anonymous):

it ask to put it in a quadratic form